# Bayesian Optimisation of Breit-Wheeler Pair Production in Simulated Laser Experiments

### v1

@@ 3.1 Simulating Shot-to-Shot Jitter
Now we have demonstrated that particle splitting gives significant reductions in runtime for the same level of accuracy, we are able to run many consecutive Monte-Carlo simulations to model the effect of jitter and capture pair production even at low laser intensity. By modelling hundreds or thousands of consecutive laser shots with the experimental parameters randomly drawn from the known distributions, we can better simulate a realistic laser experiment and estimate the total accumulated rate of pair production.
An example of such a simulated experiment is shown in Fig. 3. An initial electron beam with an energy of $7.5\,\mathrm{GeV}$ collides with an intense laser pulse, again with $25\,\mathrm{J}$ in $25\,\mathrm{fs}$ focussed to a $2\,\mathrm{\mu m}$ spot. Now the prior distributions of laser pointing $\Delta x$ and $\Delta y$ (both $\sim\mathcal{N}(0,10\,\mathrm{\mu m})$) and timing jitter $\Delta t$ $\sim\mathcal{N}(0,25\,\mathrm{fs})$ lead to imperfect collisions with a distribution of $\chi_{e}$ and $\chi_{\gamma}$. The posterior distribution for the number of pairs produced $N_{+}/N_{-}$ includes both an error from shot noise and a much more substantial variation from changes in jitter. This simulated experiment took 500 seconds to run on a single core, giving an average rate of pair production of $N_{+}/N_{-}\approx 0.005$. This is more than an order of magnitude lower than for the ideal ‘head-on’ collision with $\Delta x=\Delta y=\Delta t=0$, which gives a positron rate of $N_{+}/N_{-}\approx 0.09$.
@@ 3.2 Optimising Experimental Conditions
To begin with, we used a toy model with a single input variable of the split $s$ between energy in the colliding laser beam $E_{\mathrm{coll}}=sE_{\mathrm{tot}}$, and energy in the accelerating laser beam $(1-s)E_{\mathrm{tot}}$, which generates the electron beam. We initially considered a constant acceleration efficiency, with the electron energy described by $E_{e}\mathrm{[GeV]}=0.1(1-s)E_{\mathrm{tot}}\mathrm{[J]}$, such that $10\,\mathrm{J}$ in the accelerator beam produces a $1\,\mathrm{GeV}$ electron beam. The total laser energy available was set as $E_{\mathrm{tot}}=100\,\mathrm{J}$.
@@
The results from the GPR algorithm are shown in the top panels of Fig. 4 as more simulated experiments are run at different values of the stand-off distance $l_{0}$. Initially, after just $n=3$ initial random points, the model carries a high uncertainty across most of the domain. As more data points are taken this uncertainty gets smaller until by $n=10$ the function and the maximum are known accurately. The location for each set of simulated experiments is selected using the expected improvement acquisition function, shown in the bottom panels. The next set of simulated experiments are run where the expected improvement is high, narrowing down the possibilities. As more points are sampled and the confidence of the model increases, the expected improvement falls. Ultimately, this gives a maximum of $0.0080\pm 0.0005$ positrons per initial electron at a stand-off distance of $l_{0}=1.6\,\mathrm{mm}$.
@@
Figure 4: Top panels) Positron production at different stand-off distances $l_{0}$ predicted by simulated experiments, overlaid on the surrogate model produced by Gaussian Process Regression. Circles show the mean of $log_{10}(N_{+}/N_{-})$ while the error bars show the standard deviation over 1000 simulated shots, including jitter. The model’s predictions after each number of points $n$ are shown by the black line and the standard deviation on the model is shown by the shaded region. New points are shown in purple while previous points are shown in blue. In the final panel ($n=10$) a comparison is shown with a second optimisation, this time using a jitter of $40\,\mathrm{\mu m}$. Bottom panels) The expected improvement acquisition function predicted by the model against the stand-off distance. Peaks in the expected improvement are marked with dots and are used to calculate the next location to sample at $n+1$.
We can also look at the impact of changing the other experimental conditions, such as the jitter. Increasing the spatial jitter from a standard deviation of $10\,\mathrm{\mu m}$ to a larger jitter of $40\,\mathrm{\mu m}$ causes a reduction in the positron production rate, as more shots miss, and a change in the optimal parameters. Fig. 4 shows the result of the higher jitter in the right-most panel at $n=10$. Here the maximum rate of positron produciton is instead reached at
3.3 Optimum parameters for Breit-Wheeler pair production
@@
Now we have demonstrated a one-dimensional optimisation, we will use Gaussian Process Regression to maximise the number of positrons produced across several parameters in simulated experiments. To find the optimum parameters, we consider changes in the energy split $s$, the laser waist $w_{0}$ and duration $t_{0}$, and the stand-off distance $l_{0}$. A smaller laser spot and shorter duration means the laser reaches a higher intensity, enabling larger quantum parameters $\chi_{e}$ and $\chi_{\gamma}$. However, it also provides a smaller target such that when the spatial or temporal jitter is large, the electron beam is more likely to miss the laser pulse completely and encounter only a very low quantum parameter. Similarly, a smaller stand-off distance means the electron bunch is small at the collision point, with almost all electrons encountering a high laser intensity, but it increases the probability of the electron beam missing the colliding laser entirely. Larger stand-off distances increase the probability of a collision but reduce the number of electrons taking part in the interaction. We use a particle splitting rate of $u=1000$ to better resolve the positron production rate in lower field regions of the parameter space and reduce the error from shot noise.
@@
Figure 5: Bayesian optimisation of the number of positron-electron pairs produced as an electron beam passes through an intense laser pulse in three dimensions, shown as a series of projections, with the model shown through 2D contours and 1D lineouts. Each contour plot show the surrogate model integrated over one of the three dimensions in turn, with yellow areas representing areas where more positrons are produced, and blue representing areas where fewer are produced. Coloured points show the result of 100 simulated experiments taken at that point in parameter space, with the colour representing the total radiated energy on the same colour scale. 1D projections similarly show the number of positrons produced plotted against each of the three dimensions. The results of the simulated experiments are shown as points, with errorbars representing the standard error, and the surrogate model is shown through the solid line and shaded error bar. The surrogate model is shown along a lineout taken through the point with the maximum energy radiated, shown through dashed red lines on the contour plots.
@@
Figure 5 shows the results of the Bayesian optimisation, with the maximum number of positrons now obtained at an energy split of $s=0.4$, with a broad maximum between $0.2<s<0.7$. There is only gradual variation with the stand-off distance but the maximum is again obtained at $l_{0}=1\,\mathrm{mm}$. However, the minimum laser waist now yields the largest number of positrons, with the rate of pair production falling four orders of magnitude with a factor of 10 increase in $w_{0}$. This highlights the extremely strong dependence of Breit-Wheeler pair production on the quantum parameter $\chi_{\gamma}$. Even if it means fewer collisions encounter the peak laser intensity, pair production requires the highest possible laser intensity at focus.
@@ 3.4 Optimum parameters for measuring inverse Compton scatteriing
Figure 6: Bayesian optimisation of the energy radiated by an electron beam passing through an intense laser pulse in three dimensions, shown as a series of 2D contours and 1D projections. Each contour plot show the surrogate model integrated over one of the three dimensions in turn, with yellow areas representing areas where the total radiated energy is higher, and blue representing areas where it is lower. Coloured points show the result of 100 simulated experiments taken at that point in parameter space, with the colour representing the total radiated energy on the same colour scale. 1D projections similarly show the total radiated energy plotted against each of the three dimensions. The results of the simulated experiments are shown as points, with errorbars representing the standard error, and the surrogate model is shown through the solid line and shaded error bar. The surrogate model is shown along a lineout taken through the point with the maximum energy radiated, shown through dashed red lines on the contour plots.
Figure 6 shows the result of the three-dimensional maximisation of the radiated energy. Points selected for simulated experiments are rapidly focussed in on the most promising region, giving a more accurate surrogate model with lower error bars near the maximum. This allows us to quickly determine the best parameters at which to run experiments to produce the maximum amount of radiation from an electron beam colliding with a laser pulse. In this situation, with 100 J of laser energy available, the optimum is found with 66 J used to accelerate the electron beam to 6.6 GeV, and 3.4 J used to collide with it. This matches the naive expectation of maximising the peak quantum parameter $\chi_{e}\propto\gamma\sqrt{I_{0}}$ when $\gamma\propto(1-s)E_{\mathrm{tot}}$.
@@
On the other hand, when the laser waist is smaller, the optimum stand-off distance increases, as a larger electron beam here results in more electrons experiencing a high field. Similarly, a smaller laser waist requires less energy in the colliding beam to reach the same intensity, and so the optimum value of $s$ decreases with a smaller $w_{0}$. The surrogate model allows us to explore these couplings with a relatively small number of simulations.
@@ 4 Conclusion
This work was funded by EPSRC grant EP/V049461/1 and the United States Air Force Office of Scientific Research. The Viking cluster was used during this project, which is a high performance compute facility provided by the University of York. We are grateful for computational support from the University of York IT Services and the Research IT team. Data is available upon reasonable request and code is available at [15].

### v2

@@ 3.1 Simulating Shot-to-Shot Jitter
Now we have demonstrated that particle splitting gives significant reductions in runtime for the same level of accuracy, we are able to run many consecutive Monte-Carlo simulations to model the effect of jitter and capture pair production even at low laser intensity. By modelling hundreds or thousands of consecutive laser shots with the experimental parameters randomly drawn from the known distributions, we can better simulate a realistic laser experiment and estimate the total accumulated rate of pair production. Note, for benchmarking purposes we conduct most of these simulations with mono-energetic electrons, with zero energy spread and no jitter in the electron energy.
An example of such a simulated experiment is shown in Fig. 3. An initial electron beam with an energy of $7.5\,\mathrm{GeV}$ collides with an intense laser pulse, again with $25\,\mathrm{J}$ in $25\,\mathrm{fs}$ focussed to a $2\,\mathrm{\mu m}$ spot. The colliding laser parameters are chosen to reflect the designs of the ZEUS and CoReLS laser facilities [2, 27] (see also the APOLLON laser [28]), while these high electron energies have recently been achieved in guided laser wakefield acceleration experiments [29, 30, 31]. Now the prior distributions of laser pointing $\Delta x$ and $\Delta y$ (both $\sim\mathcal{N}(0,10\,\mathrm{\mu m})$) and timing jitter $\Delta t$ $\sim\mathcal{N}(0,25\,\mathrm{fs})$ lead to imperfect collisions with a distribution of $\chi_{e}$ and $\chi_{\gamma}$. The posterior distribution for the number of pairs produced $N_{+}/N_{-}$ includes both an error from shot noise and a much more substantial variation from changes in jitter. This simulated experiment took 500 seconds to run on a single core, giving an average rate of pair production of $N_{+}/N_{-}\approx 0.013\pm 0.001$. This is two orders of magnitude lower than for the ideal ‘head-on’ collision with $\Delta x=\Delta y=\Delta t=0$, which gives a positron rate of $N_{+}/N_{-}\approx 1.2$.
@@ 3.2 Optimising Experimental Conditions
To begin with, we optimised a simulated experiment with just a single variable, the stand-off distance $l_{0}$, for a fixed electron beam divergence of $1\,\mathrm{mrad}$ FWHM. The larger the stand-off distance, the larger the electron beam at the collision point, presenting a bigger target to hit. However, a larger electron beam also means a lower electron density and fewer electrons participating in the collision when one occurs. These simulations used $7.5\,\mathrm{GeV}$ electrons colliding with $E_{\mathrm{coll}}=25\,\mathrm{J}$ of laser energy at an angle of $\theta=15^{\circ}$, where the colliding laser pulse was focussed to $w_{0}=2\,\mathrm{\mu m}$ and had a duration of $25\,\mathrm{fs}$ FWHM, giving a peak power of around 1 PW.
@@
The results from the GPR algorithm are shown in the top panels of Fig. 4 as more simulated experiments are run at different values of the stand-off distance $l_{0}$. Initially, after just $n=3$ initial random points, the model carries a high uncertainty across most of the domain. As more data points are taken this uncertainty gets smaller until by $n=5$ the underlying function near the maximum and the position of the maximum are known accurately, while uncertainty remains elsewhere in the domain. Further points up to $n=10$ continue to reduce the error near the maximum. After each simulation, the next set of simulated experiments are run where the expected improvement (shown by the bottom panels) is high, narrowing down the possibilities. As more points are sampled and the confidence of the model increases, the expected improvement falls. With the ‘explore’ parameter set to just $0.2$, the algorithm increases the expected improvement in the region close to the current maximum rather than in other regions of the domain. Ultimately, this gives a maximum of $0.015\pm 0.001$ positrons per initial electron at a stand-off distance of $l_{0}=17\,\mathrm{mm}$.
@@
Figure 4: Top panels) Positron production at different stand-off distances $l_{0}$ predicted by simulated experiments, overlaid on the surrogate model produced by Gaussian Process Regression. Circles show the mean of $log_{10}(N_{+}/N_{-})$ while the error bars show the standard error over 100 simulated shots, including $10\,\mathrm{\mu m}$ spatial jitter and $25\,\mathrm{fs}$ temporal jitter. The model’s predictions after each number of points $n$ are shown by the black line and the standard deviation on the model is shown by the shaded region. New points are shown in purple while previous points are shown in blue. Bottom panels) The expected improvement acquisition function predicted by the model against the stand-off distance. Peaks in the expected improvement are marked with dots and are used to calculate the next location to sample at $n+1$.
We can also look at the impact of changing the other experimental conditions, such as the jitter, with examples shown in Fig. 5. With no spatial jitter the maximum rate of pair production is $0.69\pm 0.03$ pairs produced per initial electron, with zero stand-off-distance, reflecting the high electron energy and laser intensity. However, increasing the spatial jitter causes a reduction in the average positron production rate by orders of magnitude, as the likelihood of an electron hitting the most intense part of the laser at a given stand-off distance reduces as $\sigma^{-2}$ for a spatial jitter given by $\sigma$. This effect can be ameliorated by increasing the stand-off distance and hence the size of the electron beam at the collision point, with the optimal stand-off distance increasing to $69\,\mathrm{mm}$ at $\sigma=40\,\mathrm{\mu m}$. At these large stand-off distances, the larger electron beam is easier to hit and the increase in the number of shots that hit the electron beam outweighs the disadvantage of only a small fraction of the electron beam interacting with the laser pulse on each shot. At higher spatial jitters the optimum stand-off distance increases linearly such that the optimal size of the electron beam should be comparable to the spatial jitter.
Figure 5: a) Gaussian process optimisation of the rate of pair production at different levels of spatial jitter, from $0\,\mathrm{\mu m}$ standard deviation up to $40\,\mathrm{\mu m}$ standard deviation. 100 simulations were run at each point in parameter space, with the mean and error shown by the coloured points. In each case, the results were fitted to a Gaussian Process surrogate model, with the mean shown by the black lines and an error shown by the shaded region. b) Summary of the maximum rate of pair production (left axis, shown in blue) and the optimal stand-off distance (right axis, shown in red), plotted against the level of spatial jitter. The blue dotted trend line shows a power law fit $N_{+}/N_{-}=0.7\sigma^{-1.7}$ and the orange dashed trend line shows a linear fit $l_{0}=\sigma/0.6\,\mathrm{mrad}$ respectively, for a spatial jitter given by $\sigma$.
3.3 Optimum parameters for Non-Linear Breit-Wheeler pair production
@@
Now we have demonstrated a one-dimensional optimisation, we will use Gaussian Process Regression to maximise the number of positrons produced across several parameters in simulated experiments. To find the optimum parameters, we consider changes in the energy split $s=E_{\mathrm{coll}}/E_{\mathrm{tot}}$, for $E_{\mathrm{tot}}=100\,\mathrm{J}$. The resulting electron energy is assumed to be linear with laser energy as $E_{e}=10(1-s)\,\mathrm{GeV}$. This scaling reflects all-optical inverse Compton scattering experiments over the last 13 years at a range of facilities [32, 33, 34, 35, 36]. We also vary the laser waist $w_{0}$ between $2\,\mathrm{\mu m}$ and $20\,\mathrm{\mu m}$, and the stand-off distance $l_{0}$ between $1\,\mathrm{mm}$ and $1\,\mathrm{m}$. A smaller laser spot and shorter duration means the laser reaches a higher intensity, enabling larger quantum parameters $\chi_{e}$ and $\chi_{\gamma}$. However, it also provides a smaller target such that when the spatial or temporal jitter is large, the electron beam is more likely to miss the laser pulse completely and encounter only a very low quantum parameter. Similarly, a smaller stand-off distance means the electron bunch is small at the collision point, with almost all electrons encountering a high laser intensity, but it increases the probability of the electron beam missing the colliding laser entirely. Larger stand-off distances increase the probability of a collision but reduce the number of electrons taking part in the interaction. We use a particle splitting rate of $u=1000$ to better resolve the positron production rate in lower field regions of the parameter space and reduce the error from shot noise.
Until now, we have performed the simulations for monoenergetic photons (section 2) and monoenergetic electrons (sections 3.1 and 3.2). In practice, however, LWFA electron beams typically have high variability shot-to-shot and significant energy spread. For these simulations we model energy spread by sampling electrons from a Gaussian distribution with a standard deviation $10\%$ of the electron mean energy. Shot-to-shot variation in the electron mean energy itself is modelled by sampling from a Gaussian distribution with a mean of $10(1-s)\,\mathrm{GeV}$ and a standard deviation of $(1-s)\,\mathrm{GeV}$, where $s=E_{\mathrm{coll}}/E_{tot}$ as before. This reflects approximately $10\%$ shot-to-shot variation in electron energy and $10\%$ energy spread, as in e.g. Refs. [31, 37]. The electron beam divergence remains fixed at $1\,\mathrm{mrad}$. These parameters are representative of all-optical collision experiments, but are not tied to a particular setup or facility.
@@
Figure 6: Bayesian optimisation of the number of positron-electron pairs produced as an electron beam passes through an intense laser pulse in three dimensions. Top) A 3D scatter plot of the simulated experiments, coloured by positron yield. The results are fitted to a surrogate model, which is shown as contours, each showing the model predictions averaged over one axis. Below) The results of simulated experiments are plotted against each axis in turn and compared to the surrogate model mean (red line) and $2\sigma$ error (shaded area). The model predictions are calculated along lines through the optimum experimental parameters (shown as red dashed lines in the scatter plot). The transparency of the data points reflects the distance from this line, with lighter points further away from the optimum in parameter space. Error bars on the data points show the standard deviation over the 100 simulated experiments.
@@
Figure 6 shows the results of the simulations, and the Bayesian optimisation as a series of projections through the data. The maximum number of positrons obtained in these simulated experiments with $10\,\mathrm{\mu m}$ spatial jitter is $(1.57\pm 0.18)\times 10^{-2}$}, or around 2 pairs per 100 electrons. This optimum is obtained at an energy split of $s=0.43$, giving a mean electron energy of $5.7\,\mathrm{GeV}$ colliding with $1.7\,\mathrm{PW}$ of laser power. As in the previous section, the positron yield is maximised at a stand-off distance where the electron beam size is comparable to the $10\,\mathrm{\mu m}$ spatial jitter, here at $l_{0}=22\,\mathrm{mm}$. The minimum laser waist of $2\,\mathrm{\mu m}$ gives a peak field strength of $a_{0}=132$ and yields the largest number of positrons, with the rate of pair production falling four orders of magnitude with a factor of $10$ increase in $w_{0}$. This highlights the extremely strong dependence of Breit-Wheeler pair production on the quantum parameter, which here would nominally reach up to $\chi_{e}\approx 9$ for an ideal collision with no radiation reaction. These simulated experiments demonstrate that pair production requires the highest possible laser intensity at focus, even if it means accepting a relatively large stand-off distance, with few electrons reaching the peak field strength.
@@ 3.4 Optimum parameters for measuring inverse Compton scatteriing
Figure 7: The same plotting procedure as Fig. 6 is now used to show the Bayesian optimisation of the radiation yield. Again, data points show the results of 100 simulated experiments taken at given points in parameter space, with the colour describing the radiation yield. The contours (top) show projections of the surrogate model averaged over each axis in turn, while the solid red lines and shaded regions (bottom) show the surrogate model mean and $2\sigma$ error taken along lineouts through the optimum point in parameter space. The error bars on the data points here describe the standard error on the mean radiation yield.
Figure 7 shows the result of the three-dimensional maximisation of the radiated energy. Points selected for simulated experiments are rapidly focussed in on the most promising region, giving a more accurate surrogate model with lower error bars near the maximum. This allows us to quickly determine the best parameters at which to run experiments to produce the maximum amount of radiation from an electron beam colliding with a laser pulse. In this situation, with 100 J of laser energy available, the optimum is found with 68 J used to accelerate the electron beam to 6.8 GeV, and 32 J used to collide with it. This matches the naive expectation of maximising the peak quantum parameter $\chi_{e}\propto\gamma\sqrt{I_{0}}\propto(1-s)\sqrt{s}$.
@@ 4 Conclusion
This work was funded by EPSRC grant EP/V049461/1 and in part by the Air Force Office of Scientific Research (FA9550-24-1-0053). The Viking cluster was used during this project, which is a high performance compute facility provided by the University of York. We are grateful for computational support from the University of York IT Services and the Research IT team. Data is available upon reasonable request and code is available at [38].